Find the roots of the following quadratic equations by factorisation: 
2x^2 – x +1/8 = 0 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

2x^2 – x +1/8 = 0
Taking LHS, =1/8 (16x^2 - 8x + 1)
= 1/8 (16x^2 - 4x -4x + 1)
= 1/8 (4x(4x - 1) -1(4x - 1))
= 1/8 (4x - 1)^2
The roots of this equation, 2x^2 – x + 1/8 = 0,
are the values of x for which (4x - 1)^2= 0
Therefore,
(4x - 1) = 0 or (4x - 1) = 0
⇒ x = 1/4 or x = 1/4
Thanks